# MiCEval: Unveiling Multimodal Chain of Thought’s Quality via Image Description and Reasoning Steps

### v1

@@ Abstract
Xiongtao Zhou [^1] footnotemark: 1  Jie He2thanks:   Equal Contribution.  Lanyu Chen2  Jingyu Li4  Haojing Chen4 Víctor Gutiérrez-Basulto3Jeff Z. Pan [^2] footnotemark: 2 and   Hanjie Chen5thanks:   Corresponding author 1 Waseda University, Japan 2 School of Informatics, University of Edinburgh, UK 3 School of Computer Science and Informatics, Cardiff University, UK 4 Independent Researcher 5 Rice University alenai.tao@ruri.waseda.jp, j.he@ed.ac.uk j.z.pan@ed.ac.uk, hanjie@rice.edu
@@ 1 Introduction
Multimodal chain-of-thought (MCoT) enhances the performance of multimodal large language models (MLLMs) on complex reasoning tasks by generating explicit multi-step reasoning chains to achieve the final goal Zhang et al. (2024); Gao et al. (2023); Chen et al. (2024d); Shao et al. (2024). This approach not only improves the interpretability of MLLM outputs but also offers a valuable framework for assessing their reasoning capabilities. Most existing evaluations Lu et al. (2022); Zhang et al. (2024); Chen et al. (2024d) focus exclusively on the correctness of the final prediction in MCoT, while overlooking the quality of the reasoning process within the chains. This approach, which treats the reasoning chain as a whole, fails to capture errors and unreliable steps that may occur throughout the chain Lyu et al. (2023); Turpin et al. (2023). To establish a principled framework that more accurately reflects the quality of MCoT, it is essential to evaluate each step in the reasoning chain individually.
@@
A common approach for evaluating (unimodal) CoT is to compare model-generated reasoning chains with human-written reference chains Clinciu et al. (2021); Welleck et al. (2022); Saparov and He (2023). However, this method incurs substantial costs. As a result, recent research has shifted towards reference-free evaluation methods, fine-tuning models on human-annotated reasoning chain datasets Golovneva et al. (2022); Prasad et al. (2023). Yet, these methods often fall short in delivering robust evaluation metrics due to the inherent limitations of fine-tuning. The use of large language models (LLMs) as evaluators has emerged as a promising alternative He et al. (2024); Xia et al. (2024). However, note that these techniques have been developed specifically for CoT and cannot be directly applied to MCoT, as they are unable to effectively evaluate the image descriptions in MCoT. As shown in the first step of the first MCoT answer in Figure 1, the statement “In Pair 1, both magnets’ North poles face each other directly” is an image description. Methods that focus solely on the textual modality cannot properly evaluate its correctness. In the multimodal domain, several evaluation metrics have been proposed to assess the alignment between images and descriptions by calculating their cosine similarity Hessel et al. (2021); Li et al. (2023a). However, their effectiveness in evaluating MCoT remains unexplored.
The answers generated by MCoT involve both image descriptions and their corresponding reasoning steps. Therefore, for an MCoT to be correct, all steps related to both image descriptions and reasoning must be accurate, relevant, and contribute effectively to reaching the final answer. This highlights the need for a method that evaluates both the visual and textual components, providing a comprehensive assessment of the overall quality of the MCoT output. As noted in the context of CoT by Jacovi et al. (2024), the creation of a multi-step evaluation dataset is essential for accurately assessing the abilities of MLLMs as MCoT verifiers. Such a comprehensive dataset will not only enable detailed evaluation of an MCoT verifier’s effectiveness but will also offer insights into the overall quality of MCoT outputs.
In this paper, we propose MiCEval, a framework designed to evaluate MCoT by breaking down MCoT chains into description and reasoning steps, and assessing the quality of each. MiCEval decomposes MCoT verification into five core tasks across both visual and textual modalities: description correctness, image relevance, logic correctness, logic relevance, and informativeness. Additionally, MiCEval introduces a dataset comprising 903 MLLM-generated MCoT answers and 2,889 human-annotated MCoT steps, which serve as the foundation for our comprehensive step-wise evaluations of both MCoT and its verifiers.
@@ 2.2 MCoT-level and Step-level Evaluation
MCoT-level. Given the input $\mathcal{I,Q,RC}$, the MLLM generates a score between $[0,1]$ that reflects the correctness of the entire sequence $\mathcal{RC}$.
Step-level. For the given input $\mathcal{I}$, $\mathcal{Q}$, the current step $r_{i}$, and the previous steps $\mathcal{RC}_{i-1}$ = {$r_{1}$,$r_{2}$,…, $r_{i-1}$}, the MLLM outputs a score between $[0,1]$ that reflects the correctness of $r_{i}$.
@@ 2.3 Step Type
Step Type. In an MCoT sequence, the steps are generated from information extracted from both visual and textual modalities. We define steps derived solely from visual information (i.e., describing visual content) as description steps. Steps that involve information inferred beyond the visual content, incorporating the question and previous steps, are referred to as reasoning steps.
@@ 2.3.2 Reasoning Step
Informativeness. A step is considered “Informative” if it introduces new information and does not repeat or provide redundant details from previous steps Chen et al. (2023); Prasad et al. (2023). Otherwise, it is labeled as “Uninformative”.
@@ 4.2 High-Quality Annotation Protocol
It is important to note that the requirement for fine-grained and multi-task annotations necessitates that annotators have relevant expertise. Therefore, all annotators possess the necessary background knowledge related to the annotation tasks. For details on the preliminary training of annotators, please refer to the App. B.
@@ 7 Related Work
Reasoning Chain Evaluations. Chain of Thought (CoT) is an innovative method for solving reasoning tasks using large language models (LLMs) Wei et al. (2022). Kojima et al. (2022) showed that adding a simple prompt, such as “Let’s think step by step”, significantly improves LLMs performance in zero-shot settings. Jung et al. (2022) demonstrated that even erroneous reasoning chains can lead to correct final predictions, which has increasingly shifted the focus towards verifying the correctness of the entire reasoning chain Welleck et al. (2022); Chen et al. (2024b). Golovneva et al. (2022) introduced a comprehensive set of CoT evaluation metrics, validating CoT’s logical correctness, grammar, and informativeness. Prasad et al. (2023) evaluates the informativeness and logical correctness of CoTs, while Jacovi et al. (2024) focuses on verifying both attribution and logical correctness of CoT answers.

### v4

@@ Abstract
Xiongtao Zhou [^1] footnotemark: 1  Jie He2thanks:   Equal Contribution.  Lanyu Chen1  Jingyu Li4  Haojing Chen4 Víctor Gutiérrez-Basulto322footnotemark: 2Jeff Z. Pan [^2] footnotemark: 2 and   Hanjie Chen5thanks:   Corresponding author 1 Waseda University, Japan 2 School of Informatics, University of Edinburgh, UK 3 School of Computer Science and Informatics, Cardiff University, UK 4 Independent Researcher 5 Rice University, US alenai.tao@ruri.waseda.jp, j.he@ed.ac.uk j.z.pan@ed.ac.uk, hanjie@rice.edu
@@ 1 Introduction
Multimodal chain-of-thought (MCoT) enhances the performance of multimodal large language models (MLLMs) on complex reasoning tasks by generating explicit multi-step reasoning chains to achieve the final goal Zhang et al. (2024); Gao et al. (2023); Chen et al. (2024d); Shao et al. (2024). It not only improves the interpretability of MLLM outputs but also offers a valuable framework for assessing their reasoning capabilities. Most existing evaluations Lu et al. (2022); Zhang et al. (2024); Chen et al. (2024d) focus exclusively on the correctness of the final prediction in MCoT, while overlooking the quality of the reasoning process within the chains, failing to capture errors and unreliable steps that may occur throughout the chain Lyu et al. (2023); Turpin et al. (2023). To establish a principled framework to more accurately reflect the quality of MCoT, it is essential to evaluate each step in the reasoning chain individually.
@@
A common approach for evaluating (unimodal) CoT is to compare model-generated reasoning chains with human-written reference chains Clinciu et al. (2021); Welleck et al. (2022); Saparov and He (2023). However, this method incurs substantial costs. As a result, recent research has shifted towards reference-free evaluation methods, fine-tuning models on human-annotated reasoning chain datasets Golovneva et al. (2022); Prasad et al. (2023). Yet, these methods often fall short in delivering robust evaluation metrics due to the inherent limitations of fine-tuning. The use of large language models (LLMs) as evaluators has emerged as a promising alternative He et al. (2024); Xia et al. (2024) but they cannot be directly applied to MCoT, as they are unable to effectively evaluate the image descriptions in MCoT; e.g., in Figure 1, the statement “In Pair 1, both magnets’ North poles face each other directly” is an image description. Methods that focus solely on the textual modality cannot properly evaluate the correctness of image descriptions. In the multimodal domain, several evaluation metrics have been proposed to assess the alignment between images and descriptions by calculating their cosine similarity Hessel et al. (2021); Li et al. (2023a). However, their effectiveness in evaluating MCoT remains unexplored.
The answers generated by MCoT involve both image descriptions Chen et al. (2023b, c) and their corresponding reasoning steps. Therefore, for an MCoT to be correct, all steps related to both image descriptions and reasoning must be accurate, relevant, and contribute effectively to reaching the final answer. This highlights the need for a method that evaluates both the visual and textual components, providing a comprehensive assessment of the overall quality of the MCoT output. As noted in the context of CoT by Jacovi et al. (2024), the creation of a multi-step evaluation dataset is essential for accurately assessing the abilities of MLLMs as MCoT verifiers. Such a comprehensive dataset will not only enable detailed evaluation of an MCoT verifier’s effectiveness but will also offer insights into the overall quality of MCoT outputs.
In this paper, we propose MiCEval, a framework designed to evaluate MCoT by breaking down MCoT chains into description and reasoning steps, and assessing the quality of each. MiCEval decomposes MCoT verification into five core tasks across both visual and textual modalities: description correctness, image relevance, logical correctness, logical relevance, and informativeness. Additionally, MiCEval introduces a dataset comprising 903 MLLM-generated MCoT answers and 2,889 human-annotated MCoT steps, which serve as the foundation for our comprehensive step-wise evaluations of both MCoT and its verifiers.
@@ 2.2 MCoT-level and Step-level Evaluation
MCoT-level. Given the input $\mathcal{I,Q,RC}$, where the $\mathcal{RC}$ comprising {$r_{1}$,$r_{2}$,…, $r_{n}$} is evaluated as a whole. MLLM generates a score between $[0,1]$ that reflects the correctness of the entire sequence $\mathcal{RC}$.
Step-level. For the given input $\mathcal{I}$, $\mathcal{Q}$, the current step $r_{i}$, and the previous steps $\mathcal{RC}_{i-1}$ = {$r_{1}$,$r_{2}$,…, $r_{i-1}$}, only the current step $r_{i}$ is evaluated, the MLLM outputs a score between $[0,1]$ that reflects the correctness of $r_{i}$.
@@ 2.3 Step Type
In an MCoT sequence, the steps are generated from information extracted from both visual and textual modalities. We define steps derived solely from visual information (i.e., describing visual content) as description steps. Steps that involve information inferred beyond the visual content, incorporating the question and previous steps, are referred to as reasoning steps.
@@ 2.3.2 Reasoning Step
Informativeness. A step is considered “Informative” if it introduces new information and does not repeat or provide redundant details from previous steps Chen et al. (2023a); Prasad et al. (2023). Otherwise, it is labeled as “Uninformative”.
@@ 4.2 High-Quality Annotation Protocol
It is important to note that the requirement for fine-grained and multi-task annotations necessitates that annotators have relevant expertise. Therefore, all annotators possess the necessary background knowledge related to the annotation tasks.
@@ 7 Related Work
Reasoning Chain Construction and Evaluations. Chain of Thought (CoT) is an innovative method for solving reasoning tasks using large language models (LLMs) Wei et al. (2022). Kojima et al. (2022) showed that adding a simple prompt, such as “Let’s think step by step”, significantly improves LLMs performance in zero-shot settings. Jung et al. (2022) demonstrated that even erroneous reasoning chains can lead to correct final predictions, which has increasingly shifted the focus towards verifying the correctness of the entire reasoning chain Welleck et al. (2022); Chen et al. (2024b). To address the challenge of automatically generating good quality COTs, one idea is to connect explicit knowledge, such as knowledge graphs (KG) Pan et al. (2017b, a), and parametric knowledge from LLMs Pan et al. (2023). Wu et al. (2024) proposed to generate COTs from relevant knowledge graphs Huang et al. (2024). Wang et al. (2024c) proposed to use knowledge graph patterns, such as thouse widely used for query answering over KGs, to generate effective COT, without having to rely on the content of KGs. Golovneva et al. (2022) introduced a comprehensive set of CoT evaluation metrics, validating CoT’s logical correctness, grammar, and informativeness. Prasad et al. (2023) evaluates the informativeness and logical correctness of CoTs, while Jacovi et al. (2024) focuses on verifying both attribution and logical correctness of CoT answers.
Broadly speaking, reasoning is also related to text entailments Saadat-Yazdi et al. (2023), entailment graphs Hosseini et al. (2018); Li et al. (2024); Zhou et al. (2024), NL2SQL Vougiouklis et al. (2023); Shen et al. (2024); Zheng et al. (2024), as well as planning Wang et al. (2024b); Vyas et al. (2025) and API Wang et al. (2024a).
@@ 8 Conclusion
Acknowledgement
The computations detailed in this research were facilitated by the Edinburgh International Data Facility (EIDF) and the Data-Driven Innovation Programme at the University of Edinburgh.